OpenStudy (anonymous):

e^(0.15t)=25 solve for t

OpenStudy (anonymous):

is this what you do: ln (25) / 0.15

OpenStudy (anonymous):

t = ln(25)/0.15 = 21.46

OpenStudy (anonymous):

so im right?

OpenStudy (anonymous):

what abt this? 2^(t) = 20

OpenStudy (anonymous):

log (20) / log(2) ?

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

@Euler271

OpenStudy (anonymous):

Q= ((1.13)^t) / 20

OpenStudy (loser66):

too many comments. not sure what you want me to do

OpenStudy (anonymous):